A round piece of plate steel is 58 in. in diameter and has 35 - 1.5 in holes punched in it. How many in2 of material are left in the plate? Round your answer to two decimal places.
Answer by mananth(16946)

Area of circular steel plate = pi * r^2
r=29 in
A= pi* 29^2
=2640.74 in^2
area of hole
r= 0.75
area = 3.14*0.75^2
=1.77 in^2
no of holes = 35
total area of holes = 35*1.77
=61.82 in^2
=2640.74-61.82
2578.90 in^2 balance